HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

FT-SPECTROSCOPIES AND DEPERTURBATION ANALYSIS OF^{13}C^{18}O

FIRST EXPERIMENTAL EVIDENCE OF THE A^{1}Π ∼ a^{3}Π DIRECT INTERACTION

\underline{S. RYZNER}, A. STASIK, W. SZAJNA, Materials Spectroscopy Laboratory, Institute of Physics, University of Rzeszów, Poland; M. I. MALICKA, The Faculty of Mathematics and Applied Physics, Rzeszów University of Technology, Rzeszów, Poland; A. N. HEAYS, J. Heyrovsky Institute of Physical Chemistry, Prague, Czech Republic; W. UBACHS, Vrije Universiteit, Amsterdam, Netherlands; R. W. FIELD,

Massachusetts Institute of Technology, Cambridge, USA; N. DE OLIVEIRA, Synchrotron SOLEIL, St. Aubin, France; R. HAKALLA, Materials Spectroscopy

Laboratory, Institute of Physics, University of Rzeszów, Poland

The primary objective of this study was to perform a comprehensive investigation of the A^{1}Π(v = 2) level in the ^{13}C^{18}O isotopologue. To achieve this goal, two comprehensive FT techniques were employed to acquire the spectra: (i) emission spectroscopy in the VIS region using the Bruker IFS 125HR FT-spectrometer at the University of Rzeszów, and (ii) VUV absorption spectroscopy using the wave-front-division spectrometer as the end station on the DESIRS beamline at the SOLEIL synchrotron.

A deperturbation analysis was performed for the A^{1}Π(v = 2) level in^{13}C^{18}O using the PGOPHER software^{1}, based on 961 observed transitions from the A^{1}Π− X^{1}Σ^{+}(2, 0), B^{1}Σ^{+}− A^{1}Π(0, 2), C^{1}Σ^{+}− A^{1}Π(0, 2), D^{1}Δ− X^{1}Σ^{+}(3, 0), a’^{3}Σ^{+}− X^{1}Σ^{+}(12, 0), a^{3}Π− X^{1}Σ^{+}(13, 0), d^{3}Δ− X^{1}Σ^{+}(7, 0), e^{3}Σ^{−}− X^{1}Σ^{+}(4, 0) and I^{1}Σ^{−}− X^{1}Σ^{+}(3, 0) systems. This analysis yielded 49 molecular parameters of the investigated levels, 33 encompassing molecular constants, 16 interaction parameters, and 429 ro-vibronic terms. During the analysis, the extra-lines of the a^{3}Π(v = 13) level in the A^{1}Π − X^{1}Σ^{+}(2, 0), B^{1}Σ^{+} − A^{1}Π(0, 2) and C^{1}Σ^{+} − A^{1}Π(0, 2) bands of ^{13}C^{18}O have been observed. Thus, the direct interaction between the A^{1}Π and a^{3}Π states was detected and analysed for the first time.

Acknowledgments: RH thanks LASERLAB-EUROPE (grant numbers EUH2020-RIP-654148 and EC’s-SPF-284464) and RWF thanks the US National Science Foundation (grant number CHE-1800410) for support of this research. The authors are grateful to the general and technical staff of SOLEIL synchrotron for running the facility and providing beam time under project numbers 20120653 and 20160118.

1doi:10.1016/j.jqsrt.2016.04.010, C. Western, JQSRT, 186, 221-242, (2017).

p-number: p077  Submitted on Mon, 12 Jun 2023 11:31:38 +0200